# Measuring Agents in Production

### v1

@@ D.1 Interview Outline
D.2 Interviewees Demographics
To respect confidentiality agreements with case study sources, we present statistics on the sources in aggregate in Figure 24.
Appendix E Survey Questionnaire Details
@@
We crafted the questionnaire iteratively, refining it through early practitioner discussions. To facilitate broad participation, we limited the length, technical-depth, and disclosure-depth necessary to complete the questionnaire. All questions were optional and participation was entirely voluntary. Proceeding beyond certain questions required an answer or input however. For example, no participant could proceed without confirming they had read and understood Acknowledgement E.1-1. Figure 26 and Figure 27 shows the control-flows of the Core and Additional sections of the questionnaire respectively.
Futher, all questions are intended for practitioners building AI Agents. Respondents who reported making technical contributions to more than 1 system (Table Core Questions N1) were asked to focus all subsequent responses on 1 system of their choice (Acknowledgement E.1-2). Those who reported that they did not contribute to any systems they personally refer to as “AI agents” or “Agentic AI” systems were offered several options, such as commenting on terminology or their reason for starting the questionnaire, before being offered an early exit.
E.1 Survey Acknowledgements
@@
All participants were required to acknowledge the following statement.
You are invited to participate in a research study in which all data collected will be anonymized to protect your privacy. Your participation in this research is completely voluntary. The aim of the study is to understand key technical successes and challenges for deployment-track AI Agents and agentic systems, in order to steer further research and innovation in programming, runtime, and evaluation systems. We welcome as much or as little detail as you are willing to provide. Feedback on the survey/process itself is more than welcome. By proceeding with this survey, you acknowledge that you have read and understood the information provided above and consent to participate in this study. We sincerely appreciate your participation and value your contribution to this research. Thank you for your time and cooperation.
Acknowledgment 2
Further, all participants who stated they worked with more than 1 AI agent or agentic AI system were required to acknowledge the following statement before proceeding further.
To ensure consistency in your responses, please choose one agentic/assistant system to focus on throughout this survey. All your answers should relate to this same system. You may choose: * A system you are most familiar with, or * The system that is most developed among those you know — meaning closest to production with the most users. Please feel very, very welcome to submit additional survey responses for each system you are familiar with.
E.2 Survey Questions
@@
E.3 Participant Contribution Distribution
@@
| N1 | How many systems do you contribute to that you would personally describe as "AI agent(s)", "agentic", or "assistant(s)"? Example Answer: 2. (Required) |
@@
| N1.1 | Do your colleagues or stakeholders call the systems you work on agentic, and if so, would you be willing to answer additional questions about the one with which you are most familiar? Only shown if answer for N1 is |
@@
| N2 | May we contact you or your colleagues to learn more about your agentic/assistant systems? If so, please provide contact information. It will not be shared beyond the collaborators of this study or used for any other purpose. The question can be postponed after N14. |
@@
| N3 | With respect to the system you chose, are references available (code repositories, blogs, publications, training data, evaluation data, or benchmark links)? If so, please provide links. |
@@
| N4 | List as many keywords as you can think of describing the domains in which the target problem (opportunity) arises. |
@@
| N5 | Which of the following best describes the status of the agentic/assistant system on which you chose to focus your responses? |
@@
| N5.1 | Approximately how many target end users are actively using this production agentic system daily on average? Only show if N5 answered. |
@@
| N5.2 | Approximately, how many tasks from target end-users is the system processing per day on average? Only show if N5 answered. |
@@
| N6 | Which of the following best describes your primary role with respect to this agentic system? |
@@
| N7 | What are the ultimate target gains of enabling/deploying the system? Select the highest priority option(s). (Skip the question if you do not know.) |
@@
| N7.1 | Order the selected options according to their respective priority level. Only shown if N7 Answered. |
@@
Answers from N7 |
| N8 | Who (what) are the primary direct users or consumers of the agentic/assistant system? (Select one.) |
@@
| N9 | Referring to the previously selected description of direct users or consumers as "user" (human or non-human), what does the system require from users in terms of behavior(s), interaction(s), or role(s)? To the agent/assistant, the target end-user is an… |
@@
| N9.1 | Please sort the selected options with (1) as the primary intended role, (2) as a secondary role, and so on. Press and drag an option to sort. Only shown if N9 Answered. |
@@
| N10 | Using the previous definition of "user", how many steps or cycles can execute autonomously until user input is required? |
@@
| N11 | What determines how many steps or cycles can execute before a user’s input is required? |
@@
| N12 | How are each agent’s system prompts (instructions) constructed? |
@@
| N13 | What is the average (typical) estimated instruction length per agent in words or tokens? (Skip if you do not know.) |
@@
| N14 | What is the maximum allowable end-to-end result latency for the agentic/assistant system? |
@@
| N2 | May we contact you or your colleagues to learn more about your agentic/assistant systems? If so, please provide contact information. It will not be shared beyond the collaborators of this study or used for any other purpose. Only shown if question was postponed. |
@@
| O1 | Select any/all of the following methods currently integrated into your system that give you confidence your agentic/assistant system is consistently producing high quality outputs, whatever "high quality" means in your context. |
@@
Hover/click on the "i" for examples, definitions and descriptions.
@@
| O2 | Sort the following categories from most to least important for ongoing development and production deployment. Press and drag an option to sort. |
@@
Hover/click on the "i" for examples and definitions.
@@
| O3 | Have you compared your agentic/assistant solution to a non-agentic solution, which of the following statements is most accurate? |
@@
| O3.1 | If it were entirely up to you, would you choose the agentic solution over the alternatives? Only shown if answer for O3 is Yes |
@@
| O3.1.1 | You answered "Yes" to "If it were entirely up to you, would you choose the agentic solution over the alternatives?" Why, which of the following functional improvements does the new system offer compared to the previous solution? (Select all that apply.). Only shown if answer to O3.1 is Yes |
@@
| O3.1.1.1 | Order all selected options according to their respective priority level. - Improved performance or speed (Options from O3.1.1). Only shown if answer to O3.1.1 has answers |
@@
Answers from O3.1.1 |
| O4 | Thinking about the state of the system you chose to focus on, in your personal opinion, would you call it an "assistant" but not an AI Agent or agentic? |
@@
| O4.1 | Why do you refer to your system as AI agents or agentic? What makes it agentic in your opinion? Only shown if answer to O4 is Either/Both |
@@
| O4.2 | Why do you refer to your systems or an "assistant rather than an "agent" or "agentic"? What makes it an assistant rather than an agentic system in your opinion? You may go back and change your previous answer if you use any of these terms to describe your system. Only shown if answer to O4 is Assistant and NOT Agent or agentic |
@@
| O5 | What is the minimum level of expertise or training (knowledge and skills) expected from typical end-users? (Select the best option.) |
@@
| O6 | How would you rate the return on investment (ROI) of this agentic system relative to its total cost of development, operation, infrastructure and all other costs? (Please select the option that best reflects your assessment.) Only shown if answer to N5 is "In active production – Fully deployed and used by target end-users in a live operational environment" |
@@
| O7 | How much do each agent’s system prompt (instruction) lengths vary? (Skip if you do not know.) |
@@
| O8 | Compared to the target latency for the system’s result turn-around, how problematic is the actual latency? (Select one.) |
@@
| O9 | What is the maximum target latency for determining a single action, step, or response? (Select one.) |
@@
| O10 | What is the maximum number of distinct models used together to solve a single logical task? Skip if you do not know; estimates are fine. |
@@
| O11 | Are inference time scaling techniques used in your system? |
@@
| O11.1 | If inference time scaling techniques are used, approximately how many model calls are made per user query or task? Only shown if answer for O11 is Yes |
@@
| O12 | What data modalities does the system process now (today)? |
@@
| O13 | What data modalities will or should the system process in future (that it does NOT process now)? |
@@
Same as O12 |
| O13.1 | You have selected the following option(s) for the question "What data modalities will or should the system process in future (that it does NOT process now)?": What are the barriers to processing them now? Sort the options from most to least important; press and drag an option to sort. Only shown if O13 is answered. If you don’t know, skip the question. - No barriers, just a matter of development time |
@@
Answers selected from O13 |
| O14 | Which describes the data handling functions the system(s) perform? Select all that apply and use the "Other" box to detail. |
@@
| O15 | Are you using an openly (commercially or non-commerically) available agent-focused programming framework (e.g. LangChain, CrewAI, Autogen,…) to implement your system? |
@@
| O16 | From experimental observations, which of the following supports most of your assistant/agentic system functionality or design? Only shown if answer to O15 is Yes |
@@
| O17 | How long has it been since active prototyping or initial development started on this agentic/assistant system? |
@@
| O18 | How long have you been working on this agentic/assistant system? |

### v2

@@ C.3.2 Interview Outline
Appendix D Results on All Survey Data
In the main body of the paper, we focused on results filtered exclusively to deployed agents in production or pilot phases, to highlight successful real-world practices under realistic operational constraints. In this appendix, we present the corresponding results computed over [All Data]: all 306 valid survey responses, regardless of deployment stage. This expanded view includes prototype, research, and legacy systems (sunset and retired), providing a broader perspective across the full agent development lifecycle. For ease of comparison, each [All Data] figure mirrors a figure in the main text and uses the same layout and question wording. In the discussion below, we briefly describe the key patterns in the full dataset and highlight how they compare to the deployed-only subset.
D.1 RQ1 Agents Applications
Why Agents?
Figure 18 reproduces our analysis of motivations for using agents over non-agentic alternatives using all survey responses. We observe that the overall ranking of benefits is highly stable compared to deployed only agents as shown in Figure 1: increasing productivity remains the most frequently selected reason for adopting agents, followed by reducing human hours .
Application Domains.
For application domains of agents, (Figure 19) become even more diverse in the full dataset compared to deployed only agents (Figure 2). The same high-level industries i.e., finance and banking, technology, and corporate services remain prominent. However, the long tail of “other” domains grows, reflecting additional experimental and research systems in areas such as education, creative tools, and scientific workflows that have not yet reached deployment.
End users
latency requirements.
Latency requirements also similarly remain relaxed in the full dataset (Figure 20(b) compared to Figure 3). Most teams still report tolerating response times on the order of minutes, with a non-trivial fraction indicating that no explicit latency limit has been set. Compared to deployed agents only, the fraction of agents with undefined latency budgets is slightly higher in [All Data], which is consistent with prototypes and research artifacts that have not yet been hardened with production SLOs. Overall, these figures confirm that the preference for latency-relaxed, quality-focused applications is not an artifact of our deployment filter.
D.2 Models, Architectures, and Prompting
Autonomy and Architechture
Figure 21 explores RQ2, focusing on core component configurations and agent architectures across [All Data]. Compared to deployed agents (Figure 21(b)), the distribution of the number of distinct models used exhibits a heavier tail: non-deployed and research systems are more likely to combine many distinct models, leading to a higher incidence of configurations with four or more models. This pattern aligns with our qualitative observation that teams explore richer multi-model setups during early experimentation and then consolidate to a smaller, more manageable set of models as they move toward deployment.
A comparison between Figure 6(a) and Figure 21(a) reveals a clearer separation in autonomy. When we include all agents, systems allow a greater number of autonomous steps or cycles before human intervention compared to deployed agents only. Experimental and research systems are more likely to fall into the “tens of steps” and “no explicit limit” regimes, whereas production deployments concentrate in the low-step regime to control cost, latency, and failure amplification. Taken together, these results reinforce our interpretation that bounded autonomy is a deliberate design choice for production reliability, while higher autonomy is more common in exploratory settings.
Frameworks.
Figure 22(b) compares framework usage across all agents and corresponds to Figure 9(a) in the main text. The overall split between using a framework versus no framework remains nearly identical between the full dataset and the deployed-only subset, indicating that teams decide early whether to invest in a framework-based stack or implement their own orchestration. Within the “framework” group, the Other category expands slightly in [All Data], reflecting experimentation with a broader variety of less common or homegrown frameworks during research and prototyping stages, beyond the dominant framework families.
Prompting.
Figure 22(a) reproduces our analysis of prompt construction strategies using all survey responses. The results show that human-crafted prompts remain central across the full dataset: fully manual and manual+LLM strategies continue to be the dominant modes. However, when we include non-deployed agents, we observe a modest shift toward more automated prompting. Fully manual prompting is slightly more common among deployed agents (Figure 5), while the [All Data] distribution shows somewhat higher use of “fully autonomous” prompting and prompt optimizers. This suggests that automated prompt construction is currently used more as an experimental technique and is less frequently adopted in production systems, where controllability is critical.
Prompt lengths remain broadly similar between the deployed-only dataset (Figure 9(b)) and the full dataset (Figure 21(c)).
D.3 Evaluation Practices for All Agents
This section provides more details on evaluation practices, focusing on RQ3 across [All Data] . Figure 23 presents evaluation practices across all agents and mirrors Figure 7 and Figure 10 in the main text. Figure 23(a) shows distribution of comparison against non-agentic baselines. When we include prototypes and research systems, the fraction of teams that explicitly compare their agents to alternative solutions is slightly lower than in the deployed-only subset (34% in all data vs. 38.7% in deployed agents in Figure 10(a)). This suggests that some experimental agents are perhaps still in early stages where rigorous baseline comparison has not yet been prioritized, or where teams are primarily exploring feasibility rather than relative gains.
Figure 23(b) reports the distribution of different evaluation methods. The ordering of methods remains unchanged: human-in-the-loop (manual) evaluation is still the most common strategy, followed by model-based evaluation (LLM-as-a-judge). However, manual evaluation is somewhat more prevalent among deployed agents (Figure 7), whereas the [All Data] distribution shows a relatively higher share of automated methods. This is consistent with the idea that experimental and research systems may rely more on automated or lightweight checks, while production systems invest more heavily in human verification before and during deployment.
Figure 23(c) visualizes co-occurrence patterns between evaluation strategies. Human-in-the-loop evaluation remains the central hub in the evaluation graph, with high overlap with all other methods in the full dataset. At the same time, its co-occurrence with other strategies is slightly lower in [All Data] than in the deployed-only subset (Figure 10(b)), reflecting that some experimental systems use model-based or rule-based checks without consistently pairing them with human review. In contrast, deployed agents are more likely to combine automated evaluation with human verification perobably for higher assurance.
D.4 Challenges Across All Agent Deployment Stages
In this section, we focus on challenges encountered when building agent systems across different deployment stages, comparing deployed agents with non-deployed (i.e., prototype or research) agents in [All Data]. Specifically, we examine challenges related to data handling, latency, and modality support across all agents.
Figure 24 corresponds to Figure 12(a) and Figure 12(b) in the main text but reports statistics across all survey responses. Figure 24 shows distribution of data sources for agents which is remarkably stable when moving from deployed agents only ( Figure 12(a)) to [All Data] in Figure 24. This possibly indicates that the underlying data plumbing for agents is largely shared across lifecycle stages: teams tend to set up similar ingestion and handling pipelines during prototyping that then carry through to production with incremental hardening.
In addition, Figure 24(b) reports how often latency is described as a problem across all systems. The distribution changes only modestly when compared to the deployed-only subset (Figure 12(b)), and we again see that latency is not the dominant blocker for most agent deployments. This supports our broader conclusion that agents are currently concentrated in latency-relaxed settings where quality and correctness dominate over strict real-time responsiveness.
Appendix E Terminology
Appendix F Literature Review Scope and Categorization
Table 4 summarizes representative prior work referenced in our literature review, organized according to the type of evidence each body of work provides about AI agents. We group works into three broad categories based on their methodological focus and the nature of empirical grounding they offer.
Commercial/Industry. These studies consist of practitioner and industry reports that describe agentic systems deployed in organizational settings. These accounts often provide high-level insights into trends, adoption patterns, or organizational impact, but typically lack methodological transparency, reproducible evaluation protocols, or detailed technical characterization of deployed systems.
Research Surveys. This group of work includes academic survey papers that synthesize existing literature on AI agents, multi-agent systems, or related paradigms. While these surveys offer valuable conceptual frameworks and comprehensive overviews of prior research, they generally do not incorporate primary data from production deployments, nor do they empirically characterize how agents are built, evaluated, and operated in real-world settings.
Application/Demonstration. This category comprises works that introduce specific agentic systems, frameworks, or application-level demonstrations. These studies often showcase compelling capabilities or task-specific successes, but are typically limited to single systems or controlled scenarios and do not aim to provide field-level insights across organizations, deployment contexts, or operational constraints.
Together, these categories highlight a gap between existing conceptual, demonstrative, and industry-facing accounts and empirically grounded analyses of AI agents in production. Our study complements prior work by providing systematic, cross-organizational evidence on the technical methods, architectural patterns, and operational practices underlying deployed agentic systems.
| Category | References |
| Commercial/Industry | Capgemini Research Institute (2025), 1, PagerDuty (2025) |
| Research Surveys | Chandra et al. (2025), Chen et al. (2024), Cheng et al. (2024), Dam et al. (2024), Du et al. (2025), Guo et al. (2024) He et al. (2024), Krishnan (2025), Liu et al. (2025), Luo et al. (2025), Ma et al. (2025), Masterman et al. (2024), Mohammadi et al. (2025), Piccialli et al. (2025), Plaat et al. (2025), Shang et al. (2024), Sun et al. (2025), Tran et al. (2025), Wang et al. (2024b), Xi et al. (2023), Yehudai et al. (2025) |
| Application/Demonstration | Abramovich et al. (2024), Anthropic Engineering Team (2025), Block, Inc. (2025), Chennabasappa et al. (2025), Cline Bot Inc. (2025), Gemini CLI Maintainers (2025), Gottweis et al. (2025), Jha et al. (2025), Kon et al. (2025), Wang et al. (2024d), Movahed and Martin (2025), Schmucker et al. (2024), Shen et al. (2023), Park et al. (2023), Park et al. (2024), Parmar et al. (2025), Prabhakar et al. (2025), Singh et al. (2025), Teo et al. (2025), Wang et al. (2024a), Yang et al. (2024), Zhang et al. (2025) |
Appendix G Survey Questionnaire Details
@@
We crafted the questionnaire iteratively, refining it through early practitioner discussions. To facilitate broad participation, we limited the length, technical-depth, and disclosure-depth necessary to complete the questionnaire. All questions were optional and participation was entirely voluntary. Proceeding beyond certain questions required an answer or input however. For example, no participant could proceed without confirming they had read and understood the survey participation consent statement. Figure 28 and Figure 29 shows the control-flows of the Core and Additional sections of the questionnaire respectively. Further, all questions are intended for practitioners building AI Agents. Respondents who reported making technical contributions to more than 1 system (Table Core Questions N1) were asked to focus all subsequent responses on 1 system of their choice (Acknowledgment G.1). Those who reported that they did not contribute to any systems they personally refer to as “AI agents” or “Agentic AI” systems were offered several options, such as commenting on terminology or their reason for starting the questionnaire, before being offered an early exit.
G.1 Survey Acknowledgments
All participants who stated they worked with more than 1 AI agent or agentic AI system were required to acknowledge the following statement before proceeding further.
@@ G.1 Survey Acknowledgments
To ensure consistency in your responses, please choose one agentic system to focus on throughout this survey. All your answers should relate to this same system. You may choose:
A system you are most familiar with, or
The system that is most developed among those you know — meaning closest to production with the most users.
Please feel very, very welcome to submit additional survey responses for each system you are familiar with.
G.2 Survey Questions
@@
G.3 Participant Contribution Distribution
@@
| QN1 | How many systems do you contribute to that you would personally describe as ”AI agent(s)”, ”agentic”, or ”assistant(s)”? Example Answer: 2. (Required) |
@@
| QN1.1 | Do your colleagues or stakeholders call the systems you work on agentic, and if so, would you be willing to answer additional questions about the one with which you are most familiar? Only shown if answer for QN1 is |
@@
| QN2 | May we contact you to learn more about your agentic/assistant systems? If so, please provide contact information. It will not be shared beyond the collaborators of this study or used for any other purpose. The question can be postponed after QN14. |
@@
| QN3 | With respect to the system you chose, are references available (code repositories, blogs, publications, training data, evaluation data, or benchmark links)? If so, please provide links. |
@@
| QN4 | List as many keywords as you can think of describing the domains in which the target problem (opportunity) arises. |
@@
| QN5 | Which of the following best describes the status of the agentic/assistant system on which you chose to focus your responses? |
@@
| QN5.1 | Approximately how many target end users are actively using this production agentic system daily on average? Only show if QN5 answered. |
@@
| QN5.2 | Approximately, how many tasks from target end-users is the system processing per day on average? Only show if QN5 answered. |
@@
| QN6 | Which of the following best describes your primary role with respect to this agentic system? |
@@
| QN7 | What are the ultimate target gains of enabling/deploying the system? Select the highest priority option(s). (Skip the question if you do not know.) |
@@
| QN7.1 | Order the selected options according to their respective priority level. Only shown if QN7 Answered. |
@@
Answers from QN7 |
| QN8 | Who (what) are the primary direct users or consumers of the agentic/assistant system? (Select one.) |
@@
| QN9 | Referring to the previously selected description of direct users or consumers as ”user” (human or non-human), what does the system require from users in terms of behavior(s), interaction(s), or role(s)? To the agent/assistant, the target end-user is an… |
@@
| QN9.1 | Please sort the selected options with (1) as the primary intended role, (2) as a secondary role, and so on. Press and drag an option to sort. Only shown if QN9 Answered. |
@@
| QN10 | Using the previous definition of ”user”, how many steps or cycles can execute autonomously until user input is required? |
@@
| QN11 | What determines how many steps or cycles can execute before a user’s input is required? |
@@
| QN12 | How are each agent’s system prompts (instructions) constructed? |
@@
| QN13 | What is the average (typical) estimated instruction length per agent in words or tokens? (Skip if you do not know.) |
@@
| QN14 | What is the maximum allowable end-to-end result latency for the agentic/assistant system? |
@@
| QN2 | May we contact you or your colleagues to learn more about your agentic/assistant systems? If so, please provide contact information. It will not be shared beyond the collaborators of this study or used for any other purpose. Only shown if question was postponed. |
@@
| QO1 | Select any/all of the following methods currently integrated into your system that give you confidence your agentic/assistant system is consistently producing high quality outputs, whatever ”high quality” means in your context. |
@@
Hover/click on the ”i” for examples, definitions and descriptions.
@@
| QO2 | Sort the following categories from most to least important for ongoing development and production deployment. Press and drag an option to sort. |
@@
Hover/click on the ”i” for examples and definitions.
@@
| QO3 | Have you compared your agentic/assistant solution to a non-agentic solution, which of the following statements is most accurate? |
@@
| QO3.1 | If it were entirely up to you, would you choose the agentic solution over the alternatives? Only shown if answer for QO3 is Yes |
@@
| QO3.1.1 | You answered ”Yes” to ”If it were entirely up to you, would you choose the agentic solution over the alternatives?” Why, which of the following functional improvements does the new system offer compared to the previous solution? (Select all that apply.). Only shown if answer to QO3.1 is Yes |
@@
| QO3.1.1.1 | Order all selected options according to their respective priority level. - Improved performance or speed (Options from QO3.1.1). Only shown if answer to QO3.1.1 has answers |
@@
Answers from QO3.1.1 |
| QO4 | Thinking about the state of the system you chose to focus on, in your personal opinion, would you call it an ”assistant” but not an AI Agent or agentic? |
@@
| QO4.1 | Why do you refer to your system as AI agents or agentic? What makes it agentic in your opinion? Only shown if answer to QO4 is Either/Both |
@@
| QO4.2 | Why do you refer to your systems or an ”assistant rather than an ”agent” or ”agentic”? What makes it an assistant rather than an agentic system in your opinion? You may go back and change your previous answer if you use any of these terms to describe your system. Only shown if answer to QO4 is Assistant and NOT Agent or agentic |
@@
| QO5 | What is the minimum level of expertise or training (knowledge and skills) expected from typical end-users? (Select the best option.) |
@@
| QO6 | How would you rate the return on investment (ROI) of this agentic system relative to its total cost of development, operation, infrastructure and all other costs? (Please select the option that best reflects your assessment.) Only shown if answer to QN5 is ”In active production – Fully deployed and used by target end-users in a live operational environment” |
@@
| QO7 | How much do each agent’s system prompt (instruction) lengths vary? (Skip if you do not know.) |
@@
| QO8 | Compared to the target latency for the system’s result turn-around, how problematic is the actual latency? (Select one.) |
@@
| QO9 | What is the maximum target latency for determining a single action, step, or response? (Select one.) |
@@
| QO10 | What is the maximum number of distinct models used together to solve a single logical task? Skip if you do not know; estimates are fine. |
@@
| QO11 | Are inference time scaling techniques used in your system? |
@@
| QO11.1 | If inference time scaling techniques are used, approximately how many model calls are made per user query or task? Only shown if answer for QO11 is Yes |
@@
| QO12 | What data modalities does the system process now (today)? |
@@
| QO13 | What data modalities will or should the system process in future (that it does NOT process now)? |
@@
Same as QO12 |
| QO13.1 | You have selected the following option(s) for the question ”What data modalities will or should the system process in future (that it does NOT process now)?”: What are the barriers to processing them now? Sort the options from most to least important; press and drag an option to sort. Only shown if QO13 is answered. If you don’t know, skip the question. - No barriers, just a matter of development time |
@@
Answers selected from QO13 |
| QO14 | Which describes the data handling functions the system(s) perform? Select all that apply and use the ”QOther” box to detail. |
@@
| QO15 | Are you using an openly (commercially or non-commerically) available agent-focused programming framework (e.g. LangChain, CrewAI, Autogen,…) to implement your system? |
@@
| QO16 | From experimental observations, which of the following supports most of your assistant/agentic system functionality or design? Only shown if answer to QO15 is Yes |
@@
| QO17 | How long has it been since active prototyping or initial development started on this agentic/assistant system? |
@@
| QO18 | How long have you been working on this agentic/assistant system? |
@@
Analysis Aspects Overview
Given the diversity of agent applications and architectures, we pinpointed 17common aspects for analysis, summarized below. Regarding respective data collection, see the interview guide (Appendix C.3) and example survey question references (e.g., QN13, QO7 with full text in Appendix G).
1.
Prompt instruction length (averages, variance, distributional properties) – e.g., QN13, QO7
2.
Prompt instruction construction methods (automated, semi-automated, manual, hybrid) – e.g., QN12
3.
Model selection (open- versus closed, post-training)
4.
Number of models used (single-model vs. multi-model architectures) – e.g., QO10
5.
Use of inference-time techniques (e.g., RAG, caching, routing, speculative decoding, batching) – e.g., QO11, QO11.1
6.
Agent architecture: control flow, etc. (branching, looping, tool invocation, degree of autonomy) – e.g., QN9, QN10, QN11
7.
Agent dependencies (other agents, external software, human oversight, human expertise, data sources) – e.g., QN9, QN10, QN11, QO5
8.
Data sources – e.g., QO12
9.
Data modalities – e.g., QO13
10.
Programming frameworks – e.g., QO15, QO16
11.
Metrics – e.g., QN7
12.
Evaluation and verification methods (trustworthiness, accuracy, safety) – e.g., QO1, QO3
13.
Baseline evaluations / (human, agents, or non-agent systems) – e.g., QO3.1.1
14.
System throughput and latency – e.g., QN5.2, QN14, QO8, QO9
15.
Technical challenges (current limitations and ongoing development priorities) – e.g., QO2
16.
Application domain (domain(s), tasks, and deployment context) – e.g., QN4, QN3, QN7, QN8
17.
System stage (prototype, pilot, partial deployment, production) – e.g., QN3, QO17, QN5, QN5.1, QN5.2